OpenStudy (anonymous):

a key code must contain 4 letters and 3 numbers. the letter maybe any letter of the alphabet the number should be any number from 0 to 9.how many different keycode combunations are there

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

@DaBest21

OpenStudy (welshfella):

you need to find 26P4 for the letters and 10P3 for the numbers

OpenStudy (anonymous):

Well, If the numbers all go first it will be 6^3*26^4 Now if they are all mixed up you are testing me. Probability is a bit of a joke on this forum. Now let me see. there are 7 items in the code. I'm going for 7C3*6^3*26^4 =

OpenStudy (welshfella):

- these are the number of permutations

OpenStudy (welshfella):

one thing- i'm assuming that you cant have duplicate letters and numbers

OpenStudy (anonymous):

do 1/26*1/25*1/24*1/23*1/10*1/9*1/8 cross multiply that so do 1*1*1*1*1*1*1=1 and what is 26*25*24*23*10*9*8 @thisotherliz

OpenStudy (welshfella):

if repetitions are allowed then for the letters there are 26^4 permutations for the numbers its 10^3

OpenStudy (welshfella):

so total possible combinations = 26^4 * 10^3